Find using distributive property: 
504 x 35 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 504 × 35
= (500 + 4) × 35
= 500 × 35 + 4 × 35
= 17500 + 140
= 17640

Thanks